OpenStudy (jlopez):

nine fewer than half a numberis five more than four times the number. Define a variable, write an equation, and solve to find the number

OpenStudy (anonymous):

x = number 1/2x - 9 = 4x + 5 Combine like terms -9 - 5 = 4x - 1/2x -14 = 3.5x x = -4 Proof 1/2(-4) = -2 - 9 =-11 4(-4) = -16 + 5 = -11